OpenStudy (anonymous):

Determine whether each of the following is a geometric sequence, arithmetic sequence, or neither. 3, 1, –1, –3,..

OpenStudy (tkhunny):

Arithmetic has constant differences. Geometric has constant ratios. 3, 1, –1, –3,.. 1-3 = -2 -1-1 = -2 -3 - (-1) = -2 And we see a Common Difference. This is arithmetic. Knowing that it is Arithmetic, it cannot be Geometric, but let's do it, anyway. 3, 1, –1, –3,.. 1/3 = 1/3 -1/1 = -1 -3/(-1) = 3 That is NOT a constant Ratio. It is Not Geometric. If it fails both tests, it's neither.

OpenStudy (anonymous):

nice explanation @tkhunny ;) :)